# Discrete dark matter with light Dirac neutrinos

### v1

@@ VII Conclusion
We have proposed a new model of light Dirac neutrinos based on non-Abelian discrete flavour symmetry $A_{4}$ augmented with $Z_{2}\times U(1)_{L}$ where the symmetry breaking pattern leaves a remnant $\mathcal{Z}_{2}\subset A_{4}$ symmetry responsible for stabilising DM. Using the minimal flavon and additional fermion content required for realising neutrino data, we take both tree level and one-loop contributions to light Dirac neutrino masses. The remnant $\mathcal{Z}_{2}$ symmetry allows either fermion or scalar DM candidates. While scalar DM is dominantly produced thermally due to gauge and scalar portal interactions, fermion DM production relies primarily on Yukawa portal interactions. The same Yukawa portal interactions can also lead to thermal or non-thermal production of right handed neutrinos, which contribute to the effective relativistic degrees of freedom $\Delta N_{\rm eff}$. To illustrate this, we consider three different benchmark scenarios of such Yukawa couplings and perform a detailed study of neutrino mass, mixing together with DM and $\Delta N_{\rm eff}$. We also check the scope of other detection prospects like LFV decays. Table 2 summarises our results. While one of the sub-classes of Yukawa couplings, namely Case-A, allows fermion singlet DM with observable LFV rates, the corresponding $\Delta N_{\rm eff}$ remains negligible. The other two sub-classes allow scalar DM only but with observable $\Delta N_{\rm eff}$ at future CMB experiments. The LFV rates however remain tiny in these two sub-classes. Due to the rich particle spectrum in terms of several scalar doublets and singlets, the model can also have rich collider phenomenology. On the other hand, spontaneous breaking of discrete symmetries can have interesting consequences in cosmology due to the formation of domain walls. Such walls, which disappear due to soft-breaking terms, can emit stochastic gravitational waves which can be probed at different ongoing and future experiments. The same soft-breaking terms can also lead to decay of DM, keeping indirect detection prospects alive. We leave a comprehensive analysis of such additional detection prospects of our model to future works.
@@ Acknowledgements.
Appendix A $A_{4}$ Multiplication Rules:
@@
The multiplication rules of the irreducible representations are given by Altarelli and Feruglio (2010)
$\bf{1}\otimes\bf{1}=\bf{1},\bf{1^{\prime}}\otimes\bf{1^{\prime}}=\bf{1^{\prime\prime}},\bf{1^{\prime}}\otimes\bf{1^{\prime\prime}}=\bf{1},\bf{1^{\prime\prime}}\otimes\bf{1^{\prime\prime}}=\bf{1^{\prime}},\bf{3}\otimes\bf{3}=\bf{1}+\bf{1^{\prime}}+\bf{1^{\prime\prime}}+\bf{3}_{a}+\bf{3}_{s}$ (19)
where ${\bf a}$ and ${\bf s}$ in the subscript corresponds to anti-symmetric and symmetric parts respectively. Now, if we have two triplets as $A=(a_{1},a_{2},a_{3})^{T}$ and $B=(b_{1},b_{2},b_{3})^{T}$ respectively, their direct product can be decomposed into the direct sum mentioned above. The product rule for this two triplets in the $S$ diagonal basis [^2] Here $S$ is a $3\times 3$ diagonal generator of $A_{4}$. can be written as He et al. (2006)
@@
here $\omega$ ($=e^{2i\pi/3}$) is the cube root of unity.
@@ Appendix B The scalar potential
$+\lambda_{\eta 4}[\sum_{i,j}(\eta^{\dagger}\eta)_{3_{i}}(\eta^{\dagger}\eta)_{3_{j}}+{\rm h.c.}]+\lambda_{\eta H1}(\eta^{\dagger}\eta)_{1}(H^{\dagger}H)+\lambda_{\eta H2}(\eta^{\dagger}H)(H^{\dagger}\eta)$
$+[\lambda^{\prime}_{\eta H1}(\eta^{\dagger}H)(\eta^{\dagger}H)+{\rm h.c.}]+\big{[}\lambda_{\eta H3}(\eta^{\dagger}\eta)_{3_{1}}H^{\dagger}\eta+{\rm h.c.}\big{]}+\big{[}\lambda_{\eta H4}(\eta^{\dagger}\eta)_{3_{2}}H^{\dagger}\eta+{\rm h.c.}\big{]}$
$+\lambda_{\chi 1}(\chi\chi)_{1}(\chi\chi)_{1}+\lambda_{\chi 2}(\chi\chi)_{1^{\prime}}(\chi\chi)_{1^{\prime\prime}}+\lambda_{\chi 3}\big{[}\sum_{i,j}(\chi\chi)_{3_{i}}(\chi\chi)_{3_{j}}\big{]}+\lambda_{\chi H1}(\chi\chi)_{1}(H^{\dagger}H)$
$+\lambda_{\eta\chi 1}(\eta^{\dagger}\eta)_{1}(\chi\chi)_{1}+\lambda_{\eta\chi 2}(\eta^{\dagger}\eta)_{1^{\prime}}(\chi\chi)_{1^{\prime\prime}}+\lambda_{\eta\chi 3}(\eta^{\dagger}\eta)_{1^{\prime\prime}}(\chi\chi)_{1^{\prime}}+\lambda_{\eta\chi 4}\sum_{i,j}(\eta^{\dagger}\eta)_{3_{i}}(\chi\chi)_{3_{j}}$
$+\big{[}\lambda_{\eta H\chi}(\chi\chi)_{3_{1}}(H^{\dagger}\eta)+{\rm h.c.}\big{]}+\big{[}\lambda^{\prime}_{\eta H\chi}(\chi\chi)_{3_{2}}(H^{\dagger}\eta)+{\rm h.c.}\big{]}$
where, the product of two triplets contracted into one of the two triplet representations of $A_{4}$ is indicated by $[...]_{3_{1,2}}$, and the product of two triplets contracted into a singlet is represented by $[...]_{1,1^{\prime},1^{\prime\prime}}$. Among the $A_{4}$ triplet scalars $\eta=(\eta_{1},\eta_{2},\eta_{3})$ and $\chi=(\chi_{1},\chi_{2},\chi_{3})$, only $\eta_{1}$ and $\chi_{1}$ are getting VEV $v_{\eta}$ and $v_{\chi}$ respectively. The scalar fields can be can be expressed as
@@
$\eta_{1}$ $=$ $\begin{pmatrix}\eta_{1}^{\pm}\\ (v_{\eta}+h_{1}+iA_{1})/\sqrt{2}\end{pmatrix};\quad\eta_{2,3}=\begin{pmatrix}\eta_{2,3}^{\pm}\\ (h_{2,3}+iA_{2,3})/\sqrt{2}\end{pmatrix}$ (27)
$\chi_{1}$ $=$ $(v_{\chi}+x_{1})/\sqrt{2};\quad\chi_{2,3}=x_{2,3}/\sqrt{2}.$ (28)
@@
$\mu^{2}_{H}=\lambda_{H}v^{2}+(\lambda_{\eta H1}+\lambda_{\eta H2}+2\lambda^{\prime}_{\eta H1})\frac{v^{2}_{\eta}}{2}+\lambda_{\chi H1}\frac{v^{2}_{\chi}}{2},$ (29)
$\mu^{2}_{\eta}=(\lambda_{\eta 1}+\lambda_{\eta 2})v^{2}_{\eta}+(\lambda_{\eta H1}+\lambda_{\eta H2}+2\lambda^{\prime}_{\eta H1})\frac{v^{2}}{2}+(\lambda_{\eta\chi 1}+\lambda_{\eta\chi 2}+\lambda_{\eta\chi 3})\frac{v^{2}_{\chi}}{2},$ (30)
$\mu^{2}_{\chi}=(\lambda_{\chi 1}+\lambda_{\chi 1})v^{2}_{\chi}+\lambda_{\chi H1}\frac{v^{2}}{2}+(\lambda_{\eta\chi 1}+\lambda_{\eta\chi 2}+\lambda_{\eta\chi 3})\frac{v^{2}_{\eta}}{2}.$ (31)
@@
$M^{2}_{S+}=\begin{pmatrix}M^{2}_{hh}&M^{2}_{hh_{1}}&M^{2}_{hx_{1}}\\ M^{2}_{hh_{1}}&M^{2}_{h_{1}h_{1}}&M^{2}_{h_{1}x_{1}}\\ M^{2}_{hx_{1}}&M^{2}_{h_{1}x_{1}}&M^{2}_{x_{1}x_{1}}\\ \end{pmatrix}$ (32)
@@
$M^{2}_{hh}=8\lambda_{H}v^{2}+2(\lambda_{\eta H1}+\lambda_{\eta H2}+2\lambda^{\prime}_{\eta H1})v^{2}_{\eta}+2\lambda_{\chi H1}v^{2}_{\chi},$
$M^{2}_{hh_{1}}=(\lambda_{\eta H1}+\lambda_{\eta H2}+2\lambda^{\prime}_{\eta H1})v_{\eta}v,\,\,M^{2}_{hx_{1}}=\lambda_{\chi H1}v_{\chi}v,$
$M^{2}_{h_{1}h_{1}}=8(\lambda_{\eta 1}+\lambda_{\eta 2})v^{2}_{\eta}+2(\lambda_{\eta H1}+\lambda_{\eta H2}+2\lambda^{\prime}_{\eta H1})v^{2}+2(\lambda_{\eta\chi 1}+\lambda_{\eta\chi 2}+\lambda_{\eta\chi 3})v^{2}_{\chi},$
$M^{2}_{h_{1}x_{1}}=(\lambda_{\eta\chi 1}+\lambda_{\eta\chi 2}+\lambda_{\eta\chi 3})v_{\chi}v_{\eta},\,\,M^{2}_{x_{1}x_{1}}=8(\lambda_{\chi 1}+\lambda_{\chi 2})v^{2}_{\chi}+2\lambda_{\chi H1}v^{2}+2(\lambda_{\eta\chi 1}+\lambda_{\eta\chi 2}+\lambda_{\eta\chi 3})v^{2}_{\eta}.$
$M^{2}_{S-}=\begin{pmatrix}M^{2}_{h_{2}h_{2}}&M^{2}_{h_{2}h_{3}}&M^{2}_{h_{2}x_{2}}&M^{2}_{h_{2}x_{3}}\\ M^{2}_{h_{2}h_{3}}&M^{2}_{h_{3}h_{3}}&M^{2}_{h_{3}x_{2}}&M^{2}_{h_{3}x_{3}}\\ M^{2}_{h_{2}x_{2}}&M^{2}_{h_{3}x_{2}}&M^{2}_{x_{2}x_{2}}&M^{2}_{x_{2}x_{3}}\\ M^{2}_{h_{2}x_{3}}&M^{2}_{h_{3}x_{3}}&M^{2}_{x_{2}x_{3}}&M^{2}_{x_{3}x_{3}}\\ \end{pmatrix},$ (33)
@@
$M^{2}_{h_{2}h_{2}}=2(\lambda_{\eta 1}+\lambda_{\eta 2}+2\lambda_{\eta 4})v^{2}_{\eta}+(\lambda_{\eta H1}+\lambda_{\eta H2}+2\lambda^{\prime}_{\eta H1})v^{2}+(\lambda_{\eta\chi 1}+\lambda_{\eta\chi 2}+\lambda_{\eta\chi 3})v^{2}_{\eta},$
$M^{2}_{h_{2}h_{3}}=\frac{3}{2}\lambda_{\eta H3}vv_{\eta},\,\,M^{2}_{h_{2}x_{2}}=\lambda_{\eta\chi 4}v_{\eta}v_{\chi}$
and so on. This mass squared matrix can similarly be diagonalised by a $4\times 4$ orthogonal matrix involving six angles, in general, leading to physical states $\zeta_{i}$. In the scalar DM case, the lightest eigenstate of the $4\times 4$ mass matrix given above becomes the DM candidate. Similar mass matrices can be written for both $\mathcal{Z}_{2}$-odd and $\mathcal{Z}_{2}$-even pseudoscalars as well.

### v2

@@ VII Conclusion
We have proposed a new model of light Dirac neutrinos based on non-Abelian discrete flavour symmetry $A_{4}$ augmented with $Z_{2}\times U(1)_{L}$ where the symmetry breaking pattern leaves a remnant $\mathcal{Z}_{2}\subset A_{4}$ symmetry responsible for stabilising DM. Using the minimal flavon and additional fermion content required for realising neutrino data, we take both tree level and one-loop contributions to light Dirac neutrino masses. While tree level or type-I seesaw leads to a rank-1 Dirac neutrino mass matrix with only one non-zero mass eigenvalue, scotogenic or one-loop contributions can generate the other masses. Our proposal, therefore, provides the first generalisation of discrete dark matter and scoto-seesaw scenarios to light Dirac neutrinos.
The remnant $\mathcal{Z}_{2}$ symmetry allows either fermion or scalar DM candidates. While scalar DM is dominantly produced thermally due to gauge and scalar portal interactions, fermion DM production relies primarily on Yukawa portal interactions. The same Yukawa portal interactions can also lead to thermal or non-thermal production of right handed neutrinos, which contribute to the effective relativistic degrees of freedom $\Delta N_{\rm eff}$. To illustrate this, we consider three different benchmark scenarios of such Yukawa couplings and perform a detailed study of neutrino mass, mixing together with DM and $\Delta N_{\rm eff}$. We also check the scope of other detection prospects like LFV decays. Table 2 summarises our results. While one of the sub-classes of Yukawa couplings, namely Case-A, allows fermion singlet DM with observable LFV rates, the corresponding $\Delta N_{\rm eff}$ remains negligible. The other two sub-classes allow scalar DM only but with observable $\Delta N_{\rm eff}$ at future CMB experiments. The LFV rates however remain tiny in these two sub-classes. Due to the rich particle spectrum in terms of several scalar doublets and singlets, the model can also have rich collider phenomenology. On the other hand, spontaneous breaking of discrete symmetries can have interesting consequences in cosmology due to the formation of domain walls. Such walls, which disappear due to soft-breaking terms, can emit stochastic gravitational waves which can be probed at different ongoing and future experiments. The same soft-breaking terms can also lead to decay of DM, keeping indirect detection prospects alive. We leave a comprehensive analysis of such additional detection prospects of our model to future works.
@@ Acknowledgements.
Appendix A $A_{4}$ Multiplication rules
@@
The multiplication rules of the irreducible representations are given by Altarelli:2010gt
$\bf{1}\otimes\bf{1}=\bf{1},\bf{1^{\prime}}\otimes\bf{1^{\prime}}=\bf{1^{\prime\prime}},\bf{1^{\prime}}\otimes\bf{1^{\prime\prime}}=\bf{1},\bf{1^{\prime\prime}}\otimes\bf{1^{\prime\prime}}=\bf{1^{\prime}},\bf{3}\otimes\bf{3}=\bf{1}+\bf{1^{\prime}}+\bf{1^{\prime\prime}}+\bf{3}_{1}+\bf{3}_{2}$ (19)
where ${\bf a}$ and ${\bf s}$ in the subscript corresponds to anti-symmetric and symmetric parts respectively. Now, if we have two triplets as $A=(a_{1},a_{2},a_{3})^{T}$ and $B=(b_{1},b_{2},b_{3})^{T}$ respectively, their direct product can be decomposed into the direct sum mentioned above. The product rule for this two triplets in the $S$ diagonal basis [^3] Here $S$ is a $3\times 3$ diagonal generator of $A_{4}$. can be written as He:2006dk ; Memenga:2013vc ; Ishimori:2010au
@@
where $\omega$ ($=e^{2i\pi/3}$) is the cube root of unity.
@@ Appendix B The scalar potential
$+\lambda_{\eta 3}(\eta^{\dagger}\eta)_{3_{1}}(\eta^{\dagger}\eta)_{3_{1}}+\lambda_{\eta 4}[(\eta^{\dagger}\eta)_{3_{1}}(\eta^{\dagger}\eta)_{3_{2}}+{\rm h.c.}]+\lambda_{\eta 5}[(\eta^{\dagger}\eta)_{3_{2}}(\eta^{\dagger}\eta)_{3_{2}}+{\rm h.c.}]$
$+\lambda_{\eta H1}(\eta^{\dagger}\eta)_{1}(H^{\dagger}H)+\lambda_{\eta H2}(\eta^{\dagger}H)(H^{\dagger}\eta)+\lambda_{\eta H3}[(\eta^{\dagger}H)(\eta^{\dagger}H)+{\rm h.c.}]$
$+\lambda_{\eta H4}\big{[}(\eta^{\dagger}\eta)_{3_{1}}H^{\dagger}\eta+{\rm h.c.}\big{]}+\lambda_{\eta H5}\big{[}(\eta^{\dagger}\eta)_{3_{2}}H^{\dagger}\eta+{\rm h.c.}\big{]}+\lambda_{\chi 1}(\chi\chi)_{1}(\chi\chi)_{1}$
$+\lambda_{\chi 2}(\chi\chi)_{1^{\prime}}(\chi\chi)_{1^{\prime\prime}}+\lambda_{\chi 3}(\chi\chi)_{3_{1}}(\chi\chi)_{3_{1}}+\lambda_{\chi H1}(\chi\chi)_{1}(H^{\dagger}H)+\lambda_{\eta\chi 1}(\eta^{\dagger}\eta)_{1}(\chi\chi)_{1}$
$+\lambda_{\eta\chi 2}\big{[}(\eta^{\dagger}\eta)_{1^{\prime}}(\chi\chi)_{1^{\prime\prime}}+{\rm h.c.}\big{]}+\lambda_{\eta\chi 3}\big{[}(\eta^{\dagger}\eta)_{1^{\prime\prime}}(\chi\chi)_{1^{\prime}}+{\rm h.c.}\big{]}+\lambda_{\eta\chi 4}(\eta^{\dagger}\eta)_{3_{1}}(\chi\chi)_{3_{1}}$
$+\lambda_{\eta\chi 5}\big{[}(\eta^{\dagger}\eta)_{3_{2}}(\chi\chi)_{3_{1}}+{\rm h.c.}\big{]}+\lambda_{\eta H\chi}\big{[}(\chi\chi)_{3_{1}}(H^{\dagger}\eta)+{\rm h.c.}\big{]}$
where, the product of two triplets contracted into one of the two triplet representations of $A_{4}$ is indicated by $[...]_{3_{1,2}}$, and the product of two triplets contracted into a singlet is represented by $[...]_{1,1^{\prime},1^{\prime\prime}}$. Among the $A_{4}$ triplet scalars $\eta=(\eta_{1},\eta_{2},\eta_{3})$ and $\chi=(\chi_{1},\chi_{2},\chi_{3})$, only $\eta_{1}$ and $\chi_{1}$ are getting VEVs $v_{\eta}$ and $v_{\chi}$ respectively. The scalar fields can be can be expressed as
@@
$\eta_{1}$ $=$ $\begin{pmatrix}\eta_{1}^{\pm}\\ (v_{\eta}+h_{1}+iA_{1})/\sqrt{2}\end{pmatrix},\quad\eta_{2,3}=\begin{pmatrix}\eta_{2,3}^{\pm}\\ (h_{2,3}+iA_{2,3})/\sqrt{2}\end{pmatrix};$ (27)
$\chi_{1}$ $=$ $(v_{\chi}+x_{1})/\sqrt{2},\quad\chi_{2,3}=x_{2,3}/\sqrt{2}.$ (28)
@@
$\mu^{2}_{H}=$ $\lambda_{H}v^{2}+(\lambda_{\eta H1}+\lambda_{\eta H2}+2\lambda_{\eta H3})\frac{v^{2}_{\eta}}{2}+\lambda_{\chi H1}\frac{v^{2}_{\chi}}{2},$
$\mu^{2}_{\eta}=$ $(\lambda_{\eta 1}+\lambda_{\eta 2})v^{2}_{\eta}+(\lambda_{\eta H1}+\lambda_{\eta H2}+2\lambda_{\eta H3})\frac{v^{2}}{2}+(\lambda_{\eta\chi 1}+2(\lambda_{\eta\chi 2}+\lambda_{\eta\chi 3}))\frac{v^{2}_{\chi}}{2},$ (29)
$\mu^{2}_{\chi}=$ $(\lambda_{\chi 1}+\lambda_{\chi 2})v^{2}_{\chi}+\lambda_{\chi H1}\frac{v^{2}}{2}+(\lambda_{\eta\chi 1}+2(\lambda_{\eta\chi 2}+\lambda_{\eta\chi 3}))\frac{v^{2}_{\eta}}{2}.$
@@
$\mathcal{M}^{2}_{S+}=\begin{pmatrix}M^{2}_{hh}&M^{2}_{hh_{1}}&M^{2}_{hx_{1}}\\ M^{2}_{hh_{1}}&M^{2}_{h_{1}h_{1}}&M^{2}_{h_{1}x_{1}}\\ M^{2}_{hx_{1}}&M^{2}_{h_{1}x_{1}}&M^{2}_{x_{1}x_{1}}\\ \end{pmatrix}_{3\times 3}$ (30)
where,
$M^{2}_{hh}=2\lambda_{H}v^{2},\,\,M^{2}_{hh_{1}}=(\lambda_{\eta H1}+\lambda_{\eta H2}+2\lambda_{\eta H3})v_{\eta}v,\,\,M^{2}_{hx_{1}}=\lambda_{\chi H1}v_{\chi}v,$
$M^{2}_{h_{1}h_{1}}=2(\lambda_{\eta 1}+\lambda_{\eta 2})v^{2}_{\eta},\,\,M^{2}_{h_{1}x_{1}}=(\lambda_{\eta\chi 1}+2(\lambda_{\eta\chi 2}+\lambda_{\eta\chi 3}))v_{\chi}v_{\eta},\,\,M^{2}_{x_{1}x_{1}}=2(\lambda_{\chi 1}+\lambda_{\chi 2})v^{2}_{\chi}.$
$\mathcal{M}^{2}_{S-}=\begin{pmatrix}M^{2}_{h_{2}h_{2}}&M^{2}_{h_{2}h_{3}}&M^{2}_{h_{2}x_{2}}&M^{2}_{h_{2}x_{3}}\\ M^{2}_{h_{2}h_{3}}&M^{2}_{h_{3}h_{3}}&M^{2}_{h_{3}x_{2}}&M^{2}_{h_{3}x_{3}}\\ M^{2}_{h_{2}x_{2}}&M^{2}_{h_{3}x_{2}}&M^{2}_{x_{2}x_{2}}&M^{2}_{x_{2}x_{3}}\\ M^{2}_{h_{2}x_{3}}&M^{2}_{h_{3}x_{3}}&M^{2}_{x_{2}x_{3}}&M^{2}_{x_{3}x_{3}}\\ \end{pmatrix}_{4\times 4},$ (31)
@@
$M^{2}_{h_{2}h_{2}}=$ $M_{h_{3}h_{3}}^{2}=\frac{1}{2}((-3\lambda_{\eta 2}+4\lambda_{\eta 3})v^{2}_{\eta}-3(\lambda_{\eta\chi 2}+\lambda_{\eta\chi 3})v_{\chi}^{2}),$
$M^{2}_{x_{2}x_{2}}=$ $M_{x_{3}x_{3}}^{2}=\frac{1}{2}[-3(\lambda_{\eta\chi 2}+\lambda_{\eta\chi 3})v_{\eta}^{2}+(-3\lambda_{\chi 2}+4\lambda_{\chi 3})v_{\chi}^{2}],$
$M^{2}_{h_{2}h_{3}}=$ $3\lambda_{\eta H4}vv_{\eta}$
$M^{2}_{h_{2}x_{2}}=$ $M_{h_{3}x_{3}}^{2}=\lambda_{\eta H\chi}vv_{\chi}+\lambda_{\eta\chi 4}v_{\eta}v_{\chi},$
$M_{h_{2}x_{3}}^{2}=$ $M_{h_{3}x_{2}}^{2}=\lambda_{\eta H\chi}vv_{\chi}$
$M_{x_{2}x_{3}}^{2}=$ $\lambda_{\eta H\chi}vv_{\eta}$
This mass squared matrix can similarly be diagonalised by a $4\times 4$ orthogonal matrix involving six angles, in general, leading to physical states $\zeta_{i}$. In the scalar DM case, the lightest eigenstate of the $4\times 4$ mass matrix given above becomes the DM candidate. Similar mass matrices can be written for both $\mathcal{Z}_{2}$-odd and $\mathcal{Z}_{2}$-even pseudoscalars as well.
The mass matrix squared for pseduo-scalars in the ($A,A_{1},A_{2},A_{3}$) basis can be written as
$\mathcal{M}_{PS}^{2}=\begin{pmatrix}M^{2}_{AA}&M^{2}_{AA_{1}}&0&0\\ M^{2}_{AA_{1}}&M^{2}_{A_{1}A_{1}}&0&0\\ 0&0&M^{2}_{A_{2}A_{2}}&M^{2}_{A_{2}A_{3}}\\ 0&0&M^{2}_{A_{2}A_{3}}&M^{2}_{A_{3}A_{3}}\end{pmatrix}_{4\times 4}$ (32)
@@
$M^{2}_{AA}=$ $-2\lambda_{\eta H3}v_{\eta}^{2},\quad M^{2}_{A_{1}A_{1}}=-2\lambda_{\eta H3}v^{2},\quad M^{2}_{AA_{1}}=2\lambda_{\eta H3}vv_{\eta},$
$M_{A_{2}A_{2}}^{2}=$ $M_{A_{3}A_{3}}^{2}=\frac{1}{2}[-4\lambda_{\eta H3}v^{2}-(3\lambda_{\eta 2}+8\lambda_{\eta 5})v_{\eta}^{2}-3(\lambda_{\eta\chi 2}+\lambda_{\eta\chi 3})v_{\chi}^{2}],$
$M^{2}_{A_{2}A_{3}}=$ $\lambda_{\eta H4}vv_{\eta}.$
Similarly for the charged scalars, the mass matrix squared in the ($h^{\pm},\eta_{1}^{\pm},\eta_{2}^{\pm},\eta_{3}^{\pm}$) basis can be expressed as
$\mathcal{M}^{2}_{CS}=\begin{pmatrix}M^{2}_{h^{\pm}h^{\mp}}&M^{2}_{h^{\pm}\eta_{1}^{\mp}}&0&0\\ M^{2}_{h^{\pm}\eta_{1}^{\mp}}&M^{2}_{\eta_{1}^{\pm}\eta_{1}^{\mp}}&0&0\\ 0&0&M^{2}_{\eta_{2}^{\pm}\eta_{2}^{\mp}}&M^{2}_{\eta_{2}^{\pm}\eta_{3}^{\mp}}\\ 0&0&M^{2}_{\eta_{2}^{\pm}\eta_{3}^{\mp}}&M^{2}_{\eta_{3}^{\pm}\eta_{3}^{\mp}}\end{pmatrix}_{4\times 4}$ (33)
$M^{2}_{h^{\pm}h^{\mp}}=$ $-(\lambda_{\eta H2}+2\lambda_{\eta H3})\frac{v_{\eta}^{2}}{2},\quad M^{2}_{\eta_{1}^{\pm}\eta_{1}^{\mp}}=-(\lambda_{\eta H2}+2\lambda_{\eta H3})\frac{v^{2}}{2},$
$M_{\eta_{2}^{\pm}\eta_{2}^{\mp}}^{2}=$ $M^{2}_{\eta_{3}^{\pm}\eta_{3}^{\mp}}=-(\lambda_{\eta H2}+2\lambda_{\eta H3})\frac{v^{2}}{2}-3\lambda_{\eta 2}\frac{v_{\eta}^{2}}{2}-3(\lambda_{\eta\chi 2}+\lambda_{\eta\chi 3})\frac{v_{\chi}^{2}}{2},$
$M^{2}_{h^{\pm}\eta_{1}^{\mp}}=$ $(\lambda_{\eta H2}+2\lambda_{\eta H3})\frac{vv_{\eta}}{2},\quad M^{2}_{\eta_{2}^{\pm}\eta_{3}^{\mp}}=\lambda_{\eta H4}vv_{\eta}.$
Clearly, the $\mathcal{Z}_{2}$-even $2\times 2$ blocks of $\mathcal{M}^{2}_{PS}$ and $\mathcal{M}^{2}_{CS}$ have rank 1 indicating the massless Goldstone modes.